### v2

# The anonymization problem in social networks

Abstract
In this paper we introduce a general version of the anonymization problem in social networks, in which the goal is to maximize the number of anonymous nodes by altering a given graph. We define three variants of this optimization problem being full, partial and budgeted anonymization. In each, the objective is to maximize the number of $k$-anonymous nodes, i.e., nodes for which there are at least $k-1$ equivalent nodes, according to a particular anonymity measure of structural node equivalence. We propose four new heuristic algorithms for solving the anonymization problem which we implement into a reusable computational framework. As a baseline, we use an edge sampling method introduced in previous work. Experiments on both graph models and 19 real-world network datasets result in three empirical findings. First, we demonstrate that edge deletion is the most effective graph alteration operation. Second, we compare four commonly used anonymity measures from the literature and highlight how the choice of anonymity measure has a tremendous effect on both the initial anonymity as well as the difficulty of solving the anonymization problem. Third, we find that the proposed algorithm that preferentially deletes edges with a larger effect on nodes at a structurally unique position consistently outperforms heuristics solely based on network structure. Our best performing algorithm retains on average 14 times more edges in full anonymization, and overall ensures a better trade-off between anonymity and data utility. In the budgeted variant, it achieves 4.8 times more anonymous nodes than the baseline. This work lays foundations for future development of algorithms for anonymizing social networks.
keywords:
Complex networks, $k$-Anonymity, Privacy, Algorithms
1 Introduction
The field of social network analysis, also known as network science, studies interactions between people for various applications and downstream tasks, including modeling disease spread [1], maximizing influence spread [2], spotting anomalies [3] and finding communities [4, 5]. To conduct such research, it is desirable to have large real-world networks of people that are as realistic as possible. However, various studies found that people can still be identified in these networks using structural information, even after removing unique identifiers, i.e., pseudonymization [6, 7, 8]. As a result, sharing or publishing such networks can lead to a breach in the privacy of the people represented.
To solve this problem, sometimes also referred to as identity anonymization [9], various works introduced methods to share or publish networks in a privacy aware manner including clustering [10, 11, 12, 13], differential privacy [14, 15, 16, 17, 18, 19, 20] and $k$-anonymity [21, 22, 7, 23]. Each of these approaches anonymizes the data in a different way. For clustering, an intermediate version of the network is shared, where nodes are clustered together into supernodes and superedges in a way that anonymity is ensured. While this representation can preserve global network properties, the local structure can not be preserved. In turn, in differential privacy, users can run a query to obtain perturbed information about the network. After each query, progressively more noise should be added to ensure privacy. Even though several downstream network analysis tasks can be addressed using differential privacy [24, 25], there are a number of practical real-world settings in which a social network dataset must first be published or shared with a third party before it can be analyzed. This can be because two parties collaborate, for example, one collects and the other analyzes the data. Therefore, in this work, we focus on the $k$-anonymity methods, which aim to create an altered anonymized version of the original full network dataset.
In $k$-anonymity approaches, the anonymity of a network as a whole is determined by partitioning the nodes into equivalence classes based on a chosen notion for node equivalence, which we refer to as the anonymity measure. A node is $k$-anonymous if it is structurally equivalent to $k-1$ other nodes, according to the measure used for anonymity. For example, if a node is $k$-anonymous with respect to the degree measure, there are $k-1$ nodes with the same degree. The choice of measure essentially determines the attacker scenario against which one protects [26]. Examples are the degree based attack [21, 27, 28, 29, 30, 31, 9] or the neighborhood attack, accounting for the exact structure of a node’s neighborhood [7, 32, 33, 34, 35, 8]. While in this manuscript we focus on unlabeled and undirected graphs, the anonymity measures can be extended to account for additional attacker knowledge such as node labels [33, 34, 36, 37], edge labels [38] or edge weights [39].
Previous work focuses on the situation where $k=2$, where the anonymity of the network is measured as uniqueness: the fraction of nodes at a structurally unique network position [7, 26]. Given a network and an anonymity measure, the aim of network anonymization is then to decrease the uniqueness to a desired level, by means of network alterations, while ensuring that the resulting network data remains suitable for its intended purposes. Thus ultimately, in network anonymization there is a trade-off between anonymity and data utility. How utility is defined depends on the applications for which the anonymized network should be used. To measure utility, we distinguish between preserving structural network properties, and performance on downstream tasks. This distinction is similar to work in statistical disclosure control [40] that distinguishes between general utility, the distribution of the data, hence describing the data itself, and specific utility, concerning results of analyzing the data. Previous work has found that altering fewer edges leads to higher utility in the form of preservation of structural network features and centrality ranking of nodes [41]. Hence, in this paper we focus on altering an as small number of edges as possible.
The time complexity of anonymizing a network with as few alterations as possible depends on the chosen anonymity measure. For example, [9] showed that anonymization for the degree measure can be done in quadratic time in terms of the number of nodes. However,  [33], which uses as measure the exact structure of the 1-neighborhood of a considered node, showed that anonymization is NP-complete. Furthermore, the objective function in the anonymization problem is neither monotonic nor submodular, elegant properties that enabled the development of efficient methods for common network analysis problems such as influence maximization [42]. As a result, existing work on the anonymization of networks focuses on (meta)heuristic algorithms. Several of these algorithms are designed for a specific measure [21, 33], while those that are measure-agnostic typically perform alterations randomly [7, 43]. In addition, some works used genetic algorithms [35, 27]. The algorithms vary in the alteration operations used, and are sometimes tailored to a specific type of network [44].
In this paper, we introduce a general version of the anonymization problem in the form of three meaningful variants, being 1) full, 2) partial and 3) budgeted anonymization. By means of experiments, we investigate different aspects of the problem. We focus on edge deletion as alteration operation, because our empirical results show that this is the most effective operation when randomly applied to a network. We introduce four new heuristic measure-agnostic anonymization algorithms, which we compare against a baseline edge sampling algorithm [7]. We anonymize against four different anonymity measures to gain insight into the effect of the chosen measure on the anonymization process, from which we choose one measure to focus on. Then, we compare the performance of the proposed algorithms on the three variants of the problem and assess the utility of (partially) anonymized networks.
Our results show the effectiveness of our heuristic based algorithms, which target specific edges based on network structure or uniqueness of the nodes. On average, compared to the baseline, our best heuristic algorithm makes 4.8 times more nodes anonymous for the budgeted variant, and in full anonymization 13.9 times more edges are preserved. Additionally, our best heuristic algorithm preserves utility at least as well as the baseline approach, given the same number of edge deletions, and results in a better trade-off between anonymity and utility. Together with this paper, we release a computational framework for anonymizing networks called ANO-NET [^1] https://github.com/RacheldeJong/ANONET that allows researchers to implement their anonymization algorithms and apply them to the different variants of the anonymization problem. The framework is accompanied by a representative real-world benchmark dataset.
In summary, the contributions of this paper are the following:
  • We introduce a general version of the anonymization problem, together with three meaningful variants.
  • We propose four new measure-agnostic heuristic based algorithms for anonymization.
  • We investigate how anonymization methods and measures affect the anonymization process.
  • We compare the proposed algorithms on real-world networks for all three variants of the anonymization problem.
  • We assess the utility of the (partially) anonymized graphs and investigate the anonymity versus utility trade-off achieved by different algorithms.
It is worth noting that while introduced for social networks, privacy can also be important for other types of networks, such as software design networks [44] or networks of companies [45]. Hence, in our set of real-world networks, we include different categories of network datasets besides social networks.
The remainder of this paper is structured as follows. First, Section 2 summarizes related work on the problem. Section 3 covers preliminaries on networks and anonymization. In Section 4, we introduce the anonymization problem and its variants. Next, in Section 5, we outline our framework for the anonymization process and present the proposed heuristic anonymization algorithms. We discuss experiments and results in Section 6. Lastly, we give a summary and directions for future work in Section 7.
2 Related work
Works on the anonymization problem introduced in the literature often aim to anonymize all nodes in the network, the first variant of the problem. They typically focus on a specific anonymity measure, and based on this create an anonymization algorithm [21, 33]. The chosen measure also affects the complexity of the problem. For example, when the chosen measure is the degree of the nodes and all edge operations are allowed, the problem could be solved in $\mathcal{O}(n^{2})$, where $n$ is the number of nodes in the network [21]. However, for a more complex measure, such as the exact structure surrounding a node, the problem is NP-hard [32]. The search space for possible solutions is tremendously large, consisting of all subsets of existing edges, or even all subsets of possible edges, when allowing for edge addition. Therefore, exact approaches are not suitable and there is a clear need for approximation algorithms to handle larger graph inputs. Below we give an overview of related work and anonymization algorithms introduced categorized by the anonymity measure used in the work.
Degree. The most simple anonymity measure is the degree of the node [21, 27, 28, 29, 30, 31, 9]. As a result, exact anonymization algorithms have been introduced that require a minimum number of alterations [21, 31, 9]. A different approach is introduced in [27], which introduces a genetic algorithm for anonymization. Most existing work introduces an anonymization algorithm specific to the degree measure, such as the UMGA [31] or KDA algorithm [9].
Neighborhood. Anonymity measures that account for more complex information are based on the exact structure of the ego network, i.e., the 1-neighborhood [7, 35, 33], or as a parameterized version, the $d$-neighborhood structure [46]. When using this measure, the anonymization problem was shown to be NP-complete [33], making exact algorithms infeasible. Consequently, work has introduced greedy algorithms that modify node neighborhoods to resemble those of similar nodes [33]. Other methods include a genetic algorithm [35] and an edge sampling method [7].
Graph invariants. The work of [26] introduced two novel measures based on graph invariants: 1) the number of nodes and edges in the $d$-neighborhood of a considered node, and 2) the degree distribution of the node’s $d$-neighborhood.
VRQ. The notion of Vertex Refinement Queries (VRQ) was introduced in the work of Hay et al. [43, 22], accounting for the degree of nodes that are at most at distance $d$ of the considered node. To anonymize the networks, algorithms were introduced that use random rewiring [43] and alterations based on finding similar neighborhoods [47]. The work in [48] proves that anonymization using this measure is NP-complete for the subset variant of the problem, for which the aim is to anonymize a given subset of nodes.
In this paper, we contribute to the existing literature by introducing and exploring three meaningful variants of the anonymization problem. We present a general framework applicable to all three variants of the anonymization problem, and contribute to existing methods in the literature by introducing several measure-agnostic anonymization algorithms. These can be used independently of the chosen anonymity measure. Additionally, we examine the impact of different anonymity measures on the anonymization process and analyze the trade-off between anonymity and utility.
3 Preliminaries
In this section we introduce concepts and terminology used throughout this paper. This includes general concepts for networks and anonymity, the alteration operations and the used anonymity measures.
3.1 Networks
We define a network as an undirected graph $G=(V,E)$ containing a set of nodes $V$ and a set of edges $\{v,w\}\in E$ between nodes $v,w\in V$. The degree of a node equals the number of connections it has: $degree(v)=|\{w:\{v,w\}\in E\}|$. The distribution of node degrees in a real-world network typically resembles a powerlaw distribution with many low degree nodes, and few high degree nodes [49].
Nodes in networks often tend to cluster. This tendency is captured for a node $v\in V$ by the clustering coefficient which equals the number of triangles it is part of, divided by the maximum number of triangles it could have: $\frac{1}{2}degree(v)(degree(v)-1)$. The clustering coefficient of a network equals the average clustering coefficient of all nodes in the network. The distance between two nodes, $dist(v,w)$, equals the number of edges that needs to be traversed to move from node $v$ to node $w$. It follows that $dist(v,v)=0$ and if there is no path between the nodes we define $dist(v,w)=\infty$. This occurs when two nodes are in a different component. The component in the network with the largest number of nodes is called the largest connected component (LCC), also referred to as the giant component.
We define the $d$-neighborhood of a node, $N_{d}(v)=(V_{N_{d}(v)},E_{N_{d}(v)})$, as the subgraph consisting of the set of nodes that are at most at distance $d$ of the considered node $v$, and the set of all edges between these nodes. Two $d$-neighborhoods are structurally indistinguishable if they are isomorphic, i.e, there exists a bijection between the two node sets that preserves all connections. We can determine whether two $d$-neighborhoods are isomorphic by comparing their canonical labeling [50]. This is a label assigned by a function $\mathcal{C}(G)$ such that two graphs have the same label value only if they are isomorphic.
Networks often consist of communities, being groups of nodes more strongly connected with each other, than with the rest of the network. Communities, i.e., clusters of tightly connected nodes, can be identified using community detection algorithms [5]. Such algorithms assign each node to a separate community, maximizing the number of intra-community connections and minimizing inter-community connections. To account for the randomness inherently present in these heuristic algorithms, resulting in different communities for the same graph, a consensus clustering step can be applied [51].
Nodes in a network play different roles, with some nodes being in a more central position than others. Various centrality measures have been introduced, which can be used to rank nodes [52]. A commonly used measure is betweenness centrality [53], which assigns a centrality score to each node based on the fraction of shortest paths that pass through this node, as denoted in Equation 1. Here $\sigma(s,t)$ equals the number of shortest paths from node $s$ to node $t$, and $\sigma(s,t|v)$ the number of such paths that pass through node $v$.
$c_{B}(v)=\sum_{u,w\in V}\frac{\sigma(u,w|v)}{\sigma(u,w)}$ (1)
3.2 Anonymity in networks
For a graph and an anonymity measure $M$, two nodes $v,w\in V$ are equivalent if $M(v)=M(w)$. Two such nodes are said to be in the same equivalence class, denoted $eq_{M}(v)$ for a node $v$. A node is $k$-anonymous if $|eq_{M}(v)|=k$. We say that $P_{M}$ denotes the partition of nodes into equivalence classes using measure $M$. When choosing $k=2$, the anonymity of a graph can be summarized using the metric of uniqueness. This is defined as the fraction of unique nodes in the network, as shown in Equation 3.
$V_{u}=\{v\in V:|eq_{M}(v)|=1\}$ (2)
$U(G)=\frac{|V_{u}|}{|V|}$ (3)
Additionally, we denote the set of unique edges, i.e., edges for which at least one of the nodes is unique, as follows.
$E_{u}=\{\{v,w\}\in E:v\in V_{u}\vee w\in V_{u}\}$ (4)
3.2.1 Alteration operations
When anonymizing a graph, the aim is to make as many nodes as possible $k$-anonymous, given a chosen measure for anonymity, by means of graph alterations. Applying these alterations results in an anonymized graph $G^{\prime}=(V^{\prime},E^{\prime})$. We experiment with the three edge alteration operations listed below. It should be noted that in many downstream tasks where the objective is to classify a particular aspect of a node, or to cluster nodes, operations that add or delete nodes are often not desired. Therefore, we do not consider node deletion or addition. Moreover, as preliminary experiments in Section 6.2 show that edge deletion is the most effective of the three, we focus on edge deletion in the remainder of the paper.
  • Deletion: $E^{\prime}\leftarrow E\setminus\{\{v,w\}\},\ s.t.\ \{v,w\}\in E$.
  • Addition: $E^{\prime}\leftarrow E\cup\{\{v,w\}\},\ s.t.\ \{v,w\}\notin E$.
  • Rewiring: $E^{\prime}\leftarrow E\setminus\{\{v,w\},\{v^{\prime},w^{\prime}\}\}\cup\{\{v,w^{\prime}\},\{v^{\prime},w\}\}$
     s.t.  $\{v,w\},\{v^{\prime},w^{\prime}\}\in E$ and $\{v,w^{\prime}\},\{v^{\prime},w\}\notin E$.
3.2.2 Measures for k-anonymity
Various measures for $k$-anonymity have been introduced in the literature. As described in [26], the choice of measure essentially determines the attacker scenario one protects against and has an effect on the uniqueness of the network. The measures both differ in what structural information is taken into account, and how far they reach. In this paper, we use the measures listed in Table 1. We choose one measure for each category in Section 2, aiming to represent the most common measures used in the literature. For completeness, the measures are parameterized by distance $d$, which indicates up to which distance structural information is taken into account when deriving equivalence classes. In this paper, we focus on $d=1$.
During the anonymization process we can assign a state to each node equal to the value computed by the anonymity measure, as denoted by column $M(v,d)$ in Table 1. The simplest measure, degree, is the degree of the node which can be determined in constant time. More complex are count and $d$-$k$-anonymity. To determine the value for a given node $v$ for these measures, the direct neighborhood of the node needs to be extracted. For count the number of nodes and edges needs to be extracted and for $d$-$k$-anonymity the canonical labeling of the neighborhood needs to be computed, which can be computationally expensive for larger neighborhoods. Lastly, the vrq measure accounts for the degrees of neighboring nodes. Compared to count and $d$-$k$-anonymity this measure is less expensive to compute, while it accounts for structural information reaching one edge beyond the neighborhood.
Table 1: Anonymity measures and their abbreviation (left two columns), definition given a graph $G=(V,E)$, considered node $v$ and distance $d$ (third column), reach type of the measure, and node set affected when altering edge $\{v,w\}$ (last two columns).
Measure Abbreviation $M(v,d)$ Measure type $A_{M}(\{v,w\})$
Degree [21, 30, 31, 9] degree $degree(v)$ Degree $\{v,w\}$
Count [46] count $(|V_{N_{d}(v)}|,|E_{N_{d}(v)}|)$ Neighborhood $V_{N_{d}(v)}\cap V_{N_{d}(w)}$
$d$-$k$-Anonymity [35, 7, 32, 8] $d$-$k$-anonymity $\mathcal{C}(N_{d}(v))$
Vertex Refinement Query [43, 22] vrq
$\{degree(u):u\in V,$
$dist(u,v)\leq d\}$
Degree iterative $V_{N_{d}(v)}\cup V_{N_{d}(w)}$
After an edge is altered, the state of a number of nodes changes. Which nodes are affected depends on the reach of the chosen measure, i.e., up to where structural information is taken into account to compute its state [26]. We let $A_{M}(\{v,w\})$ in the rightmost column of Table 1 denote the set of nodes affected by altering edge $\{v,w\}$ using anonymity measure $M$. The reach depends on the type of measure considered, and can be categorized into degree based, neighborhood based and based on the degree of nodes that are iteratively further away. Computing the set of affected nodes is computationally least expensive for the degree and iterative degree based measures. For these measures, the affected nodes are the two nodes an edge $\{v,w\}$ connects, and the union of all nodes at distance $d$ from nodes $v,w$, respectively. Determining these sets of nodes requires a computational complexity of $\mathcal{O}(1)$ and $\mathcal{O}(|E|)$, respectively. For the neighborhood type measures, a node $u$ is affected if the edge $\{v,w\}$ occurs in $N_{d}(u)$. To determine the set of affected nodes, the intersection of the sets $V_{N_{d}}(v)$ and $V_{N_{d}}(w)$ needs to be determined, resulting in a computational complexity of $\mathcal{O}(|E|^{2})$.
4 The anonymization problem
We now define the anonymization problem and its three variants in Definition 1.
Definition 1 (Anonymization problem.).
Given a network $G$, a measure for anonymity $M$, value for $k$ and one or more allowed alteration operations:
  1. 1.
    Full anonymization: Make all nodes $k$-anonymous with as few alterations to the network as possible.
  2. 2.
    Partial anonymization: Ensure that a chosen fraction $0.0\leq\alpha\leq 1.0$ of the nodes is $k$-anonymous with as few alterations to the network as possible.
  3. 3.
    Budgeted anonymization: Alter the network with at most $B$ alterations to the network to make as many nodes as possible $k$-anonymous.
Most commonly studied is full anonymization which aims to make the entire network $k$-anonymous. However, since some nodes may require more work to anonymize, we introduce the partial variant aiming to anonymize a certain fraction of the nodes. This generalizes the subset variant proposed in [48] for which the aim is to anonymize a given subset of nodes. Additionally, the budgeted variant of the problem allows up to $B$ edge alterations. This enables the user to balance the anonymity versus utility trade-off, as generally with fewer alterations, more network utility is preserved.
In the remainder of this paper we aim to better understand the anonymization problem in its three variants under different settings. Therefore, in our experiments, we first perform two preliminary experiments showing how uniqueness changes when applying edge perturbations while using various 1) alteration operations and 2) anonymity measures. Second, we compare the performance of heuristic based anonymization algorithms on the three variants of the problem and compare how the algorithms preserve utility. For the budgeted variant, we use a budget of 5% as work on robustness of network suggests that randomly altering 5% of the edges preserves a high utility overall [41, 54]. We focus on $k=2$ and measures with $d=1$, as for $d\geq 2$ many more, if not all nodes, quickly become unique, whereas this does not per se represent a realistic attacker scenario.
5 Approach
In this section we first present the general framework for anonymization. Then, we introduce the five (one baseline, four new) measure-agnostic algorithms used in the remainder of this paper.
5.1 Anonymization
To anonymize a network, we propose the framework outlined in Algorithm 1. This general framework can be used for all variants of the anonymization problem introduced in Section 4 based on the target anonymity, that is, the desired number of nodes that is at least $k$-anonymous, and budget given as input. For the full and partial variant, the budget should equal $|E|$ and the target anonymity parameter $T$ should be set to $|V|$ and $\alpha*|V|$ respectively, where $\alpha$ equals the fraction of nodes that should be anonymous for the partial variant of the problem. For the budgeted variant, one should give the budget $B$ and a target anonymity of $|V|$. The other inputs required are the original graph $G$, the partition $P_{M}$ of nodes computed using the chosen measure (Table 1), the anonymization algorithm, and recompute gap $R$, which determines after how many edge alterations the equivalence partition should be recomputed. We discuss the use and effect of this parameter on runtime and performance in Section 5.2.
1:Input: graph $G$, partition $P_{M}$, measure $M$, value of $k$, anonymization algorithm $Anonalg$, budget $B$, target anonymity $T$, recompute gap $R$
2:$G^{\prime}\leftarrow G$, $G_{best}\leftarrow G$
3:while $B>0$ and $anonymity(G^{\prime},P_{M},k)<T$ do
4:     $B^{\prime}\leftarrow min(R,B)$
5:     $E_{A}\leftarrow Anonalg(G^{\prime},P_{M},M,k,B^{\prime})$triangleright See section 5.3
6:     $G^{\prime}\leftarrow update\_graph(G^{\prime},E_{A})$
7:     $P_{M}\leftarrow update\_P_{M}(G^{\prime},P_{M},M,E_{A})$triangleright Algorithm 2
8:     $B\leftarrow B-B^{\prime}$
9:     if $anonymity(G^{\prime},P_{M},k)>anonymity(G_{best},P_{M},k)$ then
10:         $G_{best}\leftarrow G^{\prime}$
11:     end if
12:end while
13:Return: $G_{best}$
Algorithm 1 anonymization
The algorithm works as follows. During each iteration of the while loop in lines  3 to 12, a set of edges is altered. In line 5 this set is chosen by the selected anonymization algorithm (see Section 5.3). The network is altered in line 6 and the affected part of the partition is updated accordingly in line 7. The latter is done in an efficient manner, described in Algorithm 2, to avoid recomputation of node states and anonymity for the entire graph, which is a computationally expensive part of the algorithm. The remainder of the lines are bookkeeping operations to ensure the right number of edges $B^{\prime}$ is deleted each iteration and the best result, $G_{best}$, is updated. The anonymity() function returns how many nodes are $k$-anonymous in the altered graph. When the while loop terminates, either budget $B$ is depleted or target anonymity $T$ is reached. Finally, line 13 returns the graph with the highest anonymity found by the algorithm.
5.2 Recompute gap
Altering the edges changes the network structure, which in turn influences the state of nodes, and hence their anonymity. This can affect the edges chosen by anonymization algorithm if this accounts for the structure surrounding an edge, or the uniqueness of nodes. As updating the equivalence classes and datastructures used by anonymization algorithms is computationally expensive, we allow for a trade-off between accuracy and computational cost by updating the graph and recomputing the partition after $R$ alterations, rather than after each graph alteration. A recompute gap value of 1 implies that the graph and partition are updated after each edge deletion, i.e., no recompute gap, while for $B/100$ the values are updated 100 times. Results in Appendix A show that using a recompute gap of 1 results in similar anonymization performance compared to when a larger recompute gap is used, yet resulting in substantially higher runtimes.
1:Input: graph $G$, partition $P_{M}$, Measure $M$, altered edges $E_{a}$
2:$V_{a}=\bigcup_{e\in E_{a}}A_{M}(e)$triangleright Set of affected nodes (Table 1)
3:for $v\in V_{a}$ do
4:     $state=M(v)$triangleright New node state
5:     $P_{M}=update(P_{M},v,state)$
6:end for
7:Return: $P_{M}$
Algorithm 2 update_pM
Algorithm 2 describes how the equivalence partition is updated given a set of altered edges. First, line 2 finds the set of affected nodes, i.e., the set of nodes for which the state changes cf. the $A_{M}()$ function in the rightmost column of Table 1. For these nodes the state is computed in lines 4, and the equivalence classes are updated accordingly in line 5. Lastly, the updated equivalence partition is returned.
5.3 Anonymization algorithms
In this section we describe the five heuristic anonymization algorithms. Each of them works independently from the considered anonymity measure and differs in terms of what type of information they take into account. For each heuristic, edges are selected by assigning a probability to each edge and based on these, selecting the desired number of edges. As described in Algorithm 1, the goal is to return a set of edges to be altered, or more precisely in our case; deleted, as this was found to be the most promising operation in Section 6.2 compared to edge addition and rewiring.
The edge deletion operation is likely the most effective as deleting edges decreases the number of nodes and edges in the neighborhood of nodes. This in turn results in smaller neighborhoods for which fewer network structures are possible and it is less likely to be unique, possibly making more nodes anonymous. The deletion of an edge has a larger effect on the states of the nodes it connects as this removes for both nodes one neighbor, and all triangles formed with this node from the neighborhood. For other affected nodes one edge is deleted from the neighborhood. In our heuristic algorithms, we account for these findings.
The five proposed heuristic algorithms can be categorized as follows: 1) edge sampling (baseline), 2) structure based, and 3) uniqueness based.
5.3.1 Edge sampling
The baseline algorithm, ES, has been used in previous work [7]. This algorithm randomly selects which edges to delete by assigning the same probability to each edge.
$P_{\text{{es}}}(\{v,w\})=1/|E|$ (5)
5.3.2 Structure based
These heuristics target edges based on a node’s surrounding network structure. The degree heuristic does so by assigning higher probabilities if both nodes connected by the edge have a high degree. By targeting these edges, this heuristic is likely to affect more nodes.
$P_{\text{{degree}}}(\{v,w\})=\frac{min(degree(v),degree(w))}{\sum_{\{v^{\prime},w^{\prime}\}\in E}min(degree(v^{\prime}),degree(w^{\prime}))}$ (6)
The aff heuristic computes the exact number of nodes affected by deleting an edge as defined in the rightmost column of Table 1. Computing the number of affected nodes for the count anonymity measure, which we focus on in this paper, has a computational complexity of $\mathcal{O}(|V|^{2})$, as discussed in Section 3.2.2. This makes the heuristic value more expensive to compute compared to the degree heuristic.
$P_{\text{{aff}}}(\{v,w\})=\frac{|A_{M}(\{v,w\})|}{\sum_{\{v^{\prime},w^{\prime}\}\in E}|A_{M}(\{v^{\prime},w^{\prime}\})|}$ (7)
5.3.3 Uniqueness based
Since primarily unique nodes need to be altered in order to increase anonymity, this category of algorithms aims to have a larger effect on unique nodes.
The first uniqueness based algorithm, unique, is guaranteed to change the state of unique nodes, as it only targets the set of unique edges, i.e., edges connected to at least one unique node, as denoted in Equation 4. From this set, the desired number of edges $B^{\prime}$ is chosen at random. If the set of unique edges is not large enough, the remainder of the edges is randomly selected from the set of all edges in the graph.
$P_{\text{{unique}}}(\{v,w\})=\begin{cases}1/|E_{u}|&\text{if $\{v,w\}\in E_{u}$, $|E_{u}|>B^{\prime}$}\\ 1&\text{if $\{v,w\}\in E_{u}$, $|E_{u}|\leq B^{\prime}$}\\ 1/|E-E_{u}|&\text{if $\{v,w\}\notin E_{u}$, $|E_{u}|<B^{\prime}$}\\ 0&\text{otherwise}\end{cases}$ (8)
The second uniqueness based algorithm, ua, takes into account the number of unique nodes that are affected by altering the given edge. As a result, edges affecting more unique nodes are more likely to be chosen. The factor of $\frac{1}{|E|}$ is added to avoid probabilities equal to 0. Similar to the aff heuristic, this heuristic is also expensive to compute when using the count anonymity measure considered in this paper, as the number of affected nodes need to be computed.
$P_{\text{{ua}}}(\{v,w\})=\frac{|A_{{M}}(\{v,w\})\cap V_{u}|+1/|E|}{\sum_{\{v^{\prime},w^{\prime}\}\in E}(|A_{M}(\{v^{\prime},w^{\prime}\})\cap V_{u}|+1/|E|)}$ (9)
5.4 Utility
Network utility can be defined in many ways. In this work, we choose to measure utility based on the change in topological network properties, and performance on three common downstream tasks in network analysis.
  1. 1.
    Clustering coefficient. The average clustering coefficient over all nodes. This value will likely decrease as triangles are destroyed by deleting edges.
  2. 2.
    Shortest path length. The average shortest path length $d(v,w)$ over all pairs of nodes in the same component. As edges are deleted, paths are destroyed, hence the average shortest path length increases. However, when the network starts to split into components, this value can decrease.
The considered downstream tasks are the following.
  1. 1.
    Network robustness. Change in the fraction of nodes in the largest connected component (LCC). This is expected to decrease as edges are deleted. A more robust network is more resilient under attacks [55].
  2. 2.
    Centrality. Top 100 most central nodes according to betweenness centrality before and after anonymization. If the top 100 nodes did not change the value equals 1.0, and 0.0 if there are no common nodes.
  3. 3.
    Community structure. Measured by the normalized mutual information [51] between the communities [5], before and after anonymization. A value of 1.0 implies the community structure is preserved, and 0.0 that it is completely lost.
6 Results
In this section we first summarize the experimental setup and network datasets used. Second, we continue with two exploratory experiments to compare the effectiveness of different alteration operations on anonymization of graph models, and compare the effect of using different anonymity measures on real-world data. Third, we compare the anonymization algorithms based on their effectiveness on the three variants of the problem. Fourth, we assess the utility achieved by the algorithms on the different variants, and investigate the achieved trade-off between anonymity and utility. Finally, we compare the runtimes of the algorithms on the different variants.
Table 2: Real-world networks used for experiments and their properties: the number of nodes, edges, average degree, clustering coefficient, fraction of nodes in the giant component (LCC), diameter and average distance. Results: initial uniqueness using count as anonymity measure and improvement ratio of ua compared to es. Initial uniqueness values lower than 0.05, i.e., for partial anonymous networks, are shown in italics.
Improvement UA vs. ES
$|V|$ $|E|$
Avg.
deg.
Clust.
coeff
Frac.
LCC
Dia-
meter
Avg.
dist.
$U(G)$
1.
Full
2.
Partial
3.
Bud-
geted
Radoslaw emails [56] 167 3,250 38.92 0.69 1.00 5 1.97 0.766 8.0 2.3 1.9
Primary school [57] 242 8,317 68.74 0.53 1.00 3 1.73 0.975 4.5 1.8 1.0
Moreno innov. [56] 241 923 7.66 0.31 0.49 5 2.47 0.245 6.2 1.8 2.1
Gene fusion [56] 291 279 1.92 0.00 0.08 9 3.90 0.024 4.8 1.0 5.6
Copnet calls [58] 536 621 2.32 0.25 0.65 22 7.37 0.024 5.5 1.0 8.6
Copnet sms [58] 568 697 2.45 0.22 0.80 20 7.32 0.026 2.4 1.0 3.1
Copnet FB [58] 800 6,418 16.05 0.32 1.00 7 2.98 0.488 7.6 2.1 1.4
FB Reed98 [59] 962 18,812 39.11 0.33 1.00 6 2.46 0.778 5.3 2.3 1.0
Arenas email [56] 1,133 5,451 9.62 0.25 1.00 8 3.61 0.230 8.4 1.7 2.4
Euroroads [56] 1,174 1,417 2.41 0.02 0.03 62 18.37 0.003 1.8 1.0 3.7
Air traffic control [56] 1,226 2,408 3.93 0.07 1.00 17 5.93 0.042 5.4 1.0 5.6
Network science [56] 1,461 2,742 3.75 0.88 0.00 17 5.82 0.039 10.0 1.0 44.4
FB Simmons81 [59] 1,518 32,988 43.46 0.33 0.99 7 2.57 0.785 2.4 2.2 1.4
DNC emails [56] 1,866 4,384 4.70 0.59 0.98 8 3.37 0.092 15.5 1.5 1.4
Moreno health [56] 2,539 10,455 8.24 0.15 1.00 10 4.56 0.054 6.7 1.0 3.6
US power grid [56] 3,783 14,124 7.47 0.28 1.00 10 3.57 0.008 5.2 1.0 3.3
Bitcoin alpha [59] 4,941 6,594 2.67 0.11 1.00 46 18.99 0.113 66.5 1.5 1.9
GRQC collab. [60] 5,241 14,484 5.53 0.69 0.79 17 6.05 0.054 42.2 1.4 3.1
Pajek Erdős [56] 6,927 11,850 3.42 0.40 1.00 4 3.78 0.044 55.7 1.0 4.0
FB GWU54 [59] 12,193 469,528 77.02 0.22 1.00 9 2.83 0.682 - - 1.1
Enron email [60] 36,692 183,831 10.02 0.72 0.92 13 4.03 0.071 - - 2.0
FB wall 2009 [56] 45,813 183,412 8.01 0.15 0.96 18 5.60 0.033 - - 4.6
Brightkite [59] 58,228 214,078 7.35 0.27 0.97 18 4.92 0.048 - - 2.6
6.1 Experimental setup and data
We implement the algorithms introduced in Section 5 in our reusable C++ ANO-NET framework. [^2] https://github.com/RacheldeJong/ANONET For the preliminary experiments on graph models we used python and NetworkX [61], and for utility analysis we used igraph [62].
We executed experiments on graph models and real-world networks. In each experiment, all edges are sequentially deleted unless the process exceeds the runtime limit of 30 minutes. In this case, the process continues only until the required number of edges is deleted for the budgeted variant. For the four networks that this pertains to, only results on observed runtime are reported, with the goal of assessing scalability of the anonymization algorithms. For the partial and budgeted variant, we set the target anonymity to 95% of the nodes and the budget to 5% of the edges respectively. The budget is chosen based on previous work indicating that randomly deleting 5% of the edges preserves most network properties [41, 54].
The recompute gap was chosen so that the last edges are deleted in the 100th step. Including the initial anonymity measurement this yields 101 data points for each run. To account for the nondeterminism in both the anonymization algorithms and the generation of model networks, we report averages over five runs. For results showing the anonymization process and utility analysis, we additionally report the standard deviation.
For results on utility, we generate for each run the anonymized graph for all 101 data points by deleting the selected edges, and for each compute the utility metrics summarized in Section 5.4. For results on community detection, we run the Leiden algorithm [5] 10 times for each network, followed by a consensus clustering step [51].
For the graph models, we use the Erdős-Rényi (ER) model [63], where edges are chosen randomly; the Barabási-Albert (BA) model [64], accounting for the preferential attachment mechanism that generates a powerlaw degree distribution, and the Watts-Strogatz (WS) model [65] with rewire probability 0.05, which generates clustered networks (i.e., graphs with triangles). For each model, we use graphs with 500 nodes and average degrees 2, 4, 16 and 64. Properties of the real-world network datasets used can be found in Table 2.
Experiments are conducted on a machine with 512 GB RAM, 128 AMD EPYC 7702 cores at 2.0 GHz and 256 threads. Each run uses one thread, which is not shared with other processes.
6.2 Alteration operations and anonymity
Here we study the effectiveness of deleting, adding, and rewiring edges for the purpose of anonymization. Figure 1 shows results on how repeatedly applying each operation at random affects the anonymity in graph models.
A first detail to note is that the number of unique nodes when no edges are altered, varies with the average degree. Graphs with higher average degree have a higher initial uniqueness, which corresponds to findings in previous works [7, 8].
The main conclusion from these results is that random deletion is more effective than random addition or rewiring, confirming what we theorized in Section 5.3. Logically, when edges are deleted, this results in a graph with lower density with less variety in neighborhoods, which could in turn result in more anonymous nodes. On the contrary, edge addition makes the graph more dense, resulting in larger neighborhoods, while for rewiring density remains constant. For the graph models, random addition and rewiring can even result in a higher uniqueness, which is especially clear for the BA model, and for the ER and WS models when the average degree is larger than 16. However, for the WS model, with average degree 16, rewiring performs slightly better than deletion when altering a fraction of 0.2 to 0.5 of the edges. This is likely due to the many nearly identical near-cliques in the generated networks.
While these preliminary results can not rule out that edge addition and rewiring can possibly be more effective in specific cases, and could be preferred for certain applications, we choose to focus on edge deletion in the remainder of the paper.
Figure 1: Random edge addition (green), deletion (red), and rewiring (grey) applied to, using the count measure, anonymize ER (top left), BA (top right) and WS (bottom left) graphs with $|V|=500$ and average degree $\in\{2,4,16,64\}$.
6.3 Anonymization and measures
Figure 2 shows how the uniqueness changes when deleting edges using different measures for anonymity. For readability, the results are shown for a subset of the anonymization algorithms, with one algorithm for each category in Section 5.3. We depict a selection of five networks representative of the overall observed behavior for the networks listed in Table 2. This is, from left to right in Figure 2, 1) a plateau with a steep decrease, 2) linear decrease, 3) plateau with linear decrease, 4) sigmoid-like decrease and 5) initial increase before decrease. The results for the remaining networks are included in Appendix B.
Overall, we see that the measure for anonymity used has a large influence on the results; both on the starting uniqueness with which the process starts and how fast the uniqueness decreases. The simplest measure, degree, has the lowest initial uniqueness and is the easiest to anonymize for. The measure with the furthest reach, vrq, has the highest initial uniqueness, which corresponds to the findings of [26], and is the most difficult to anonymize for, even more difficult than $d$-$k$-anonymity which accounts for exact structure of the neighborhood of nodes.
An example where different measures have the same initial uniqueness but differ in how quickly they are anonymized is the “Radoslaw emails” network (left). For this network, the uniqueness starts around 0.78 for both count and $d$-$k$-anonymity, but decreases substantially faster for count. However, making all nodes anonymous, i.e., full anonymization, requires a similar amount of edge deletions for both measures.
In the “GRQC collab” network (right), the results for the es algorithm shows an increase in uniqueness before decreasing. This is likely due to the collaboration network’s many cliques. When randomly deleting edges, these cliques are destroyed, which may make certain nodes unique. If edges are targeted more specifically, e.g., using the ua heuristic, this initial increase does not occur.
In the remainder of this paper, we focus on the measure count, as this measure is a realistic real-world scenario, has a substantial initial uniqueness compared to the other measures, yet is easier to anonymize for than, for example, the $d$-$k$-anonymity measure. Using the count measure, nodes are essentially equivalent if they have the same number of neighbors, and are part of the same number of triangles.
Figure 2: Uniqueness when using four different anonymity measures (color) and three anonymization algorithms (linestyle) on real-world networks. For each network we show the fraction of deleted edges (horizontal axis) and the attained uniqueness (vertical axis).
6.4 Full and partial network anonymization
In the following, we investigate how effective the proposed algorithms from Section 5.3 are in ensuring anonymity for 100% (full anonymization) and 95% (partial anonymization) of the nodes in the network. The top of Figure 3 shows the fraction of edges preserved $\frac{|E^{\prime}|}{|E|}$ when edges are deleted until the desired level of anonymity is reached. Here, a higher fraction indicates better results, as, in general, preserving more edges overall implies a higher utility of the resulting anonymized network. We look at utility in detail in Sections 6.6 and 6.7. For the partial variant eight of the networks, including “US Power grid” and “Euroroads” retain 100% of their edges as initially less than 5% of their nodes is unique.
The results show that for some networks, such as “Radoslaw emails” and “FB Simmons81” only a small fraction of edges can be preserved. For other networks, for which more edges can be preserved, the uniqueness based algorithm ua outperforms the other algorithms. When dividing the result for the best heuristic by the result of es, we obtain the improvement ratio reported in the three rightmost columns of Table 2. For the full and partial variant, on average 13.9 and 1.8 times more edges can be preserved.
Figure 3: Results for full and partial anonymization (top), where for a network the bar height indicates the fraction of edges that can be preserved when making all nodes anonymous, and the horizontal line indicates the best result obtained for partial anonymization. For Budgeted anonymization (bottom) each bar indicates the fraction of unique nodes anonymized with a budget of 5% of the edges. Color indicates the used anonymization algorithm.
6.5 Budgeted anonymization
We now move to results for budgeted anonymization. To account for the diverse values in starting uniqueness, as indicated in Table 2, the values on the vertical axis on the bottom of Figure 3 show the fraction of unique nodes that is anonymized, i.e., $1-\frac{U(G^{\prime})}{U(G)}$, when deleting at most 5% of the edges. A higher value thus indicates that the algorithm is more effective in anonymization within the given budget.
For some of the networks, the ua algorithm is quite effective. As can be derived from Table 2, this overall holds for networks with lower initial uniqueness and higher average distance. Experiments to confirm this are included in Appendix C. However, some of the networks, such as “Radoslaw emails” and “GRQC collab”, require more than 5% deletion to obtain a substantial decrease in uniqueness. This can also be derived from Figure 2 showing a plateau for “Radoslaw emails” and an initial increase in uniqueness for “GRQC collab”.
Overall, we see a big performance difference both between networks and anonymization algorithms. For some of the networks, such as “Moreno health”, “Euroroads”, “Copnet sms” or “US power grid”, the uniqueness based anonymization algorithm, ua, manages to anonymize many more nodes compared to the other algorithms. For all networks, this algorithm either performs similar or outperforms the other algorithms. For the “Euroroads” network, ua even succeeds to anonymize all nodes within the given budget. On average, 4.8 times more nodes are anonymized compared to the baseline es algorithm. The algorithms especially perform well for networks with a lower initial uniqueness, lower average degree and higher average distance, which is further explained in Appendix C.
6.6 Utility
In this section we investigate the utility of the anonymized networks in terms of how well various network properties and performance on downstream tasks are preserved for the different variants of the anonymization problem. Figure 4 summarizes this for the five anonymization algorithms. A property is considered preserved if the value after anonymization $\pm$ the standard deviation differs less than 5% from the original value. As discussed in Section 6.1, we exclude the four largest networks. Additionally, for partial anonymization, we exclude the eight networks which are already partially anonymous, i.e., networks with an initial uniqueness less than 5%, resulting in 11 networks for this variant.
When comparing results for the three variants of the problem, we observe that for full and partial anonymization properties are usually not preserved, except for the largest component size which the ua algorithm manages to preserve this for six networks for the full variant. This difference compared to other algorithms is likely due to the fact that this algorithm requires fewer edge deletions. Moving to the budgeted variant, we see that more often properties are preserved, and there is a small difference in results between the algorithms. Hence, focusing on the budgeted variant seems like a promising way to further investigate and compare anonymization algorithms.
Focusing on the different properties, we see that the largest component size is typically preserved best: for budgeted anonymization, ua even manages to preserve this property for all networks. This is also shown in Appendix D, which shows for each utility metric the difference to the initial value changes when deleting more edges. The average shortest path length and most central nodes according to betweenness centrality are the second and third best preserved properties. The clustering coefficient is not preserved for most networks and out of the compared algorithms unique performs best. This could be due to the fact that this algorithm does not target edges that affect many nodes, which can potentially destroy many triangles. The community structure is not preserved by any algorithms for most networks. This property seems especially challenging to preserve, as deleting the first percentage of edges already results in a large difference, despite using the consensus clustering step to obtain more stable communities.
Overall, ua best preserves the properties for all variants. Compared to the other algorithms in the budgeted variant, it best preserves most properties, and performs similar in preserving the clustering coefficient and average shortest path length.
Figure 4: For each of the three variants of the anonymization problem (columns) the color and number indicate the fraction of networks for which the network property or performance on a downstream task (vertical axis) is preserved, i.e., the change $\pm$ one standard deviation is less than 5% of the original value, using the five anonymization algorithms (horizontal axis).
6.7 Anonymity vs. utility
After comparing algorithms in terms of achieved anonymity and utility, we now investigate the trade-off between the two. Figure 5 shows for two network datasets the Pareto optimal solutions in terms of uniqueness and performance on downstream tasks found by the algorithms. Hence, the figure shows only solutions such that there are no other solutions found by the same algorithm with both a lower uniqueness and lower difference in performance on the downstream task. For readability, we included two networks that show the overall behavior found in the full set of networks for which the results can be found in Appendix E.
When comparing the algorithms, we see that ua always gives either the best trade-off between utility and anonymity, or one similar to the best algorithm. Moreover, on the downstream tasks, the ua algorithm always gives strictly better values than es, which often gives the worst trade-off of all algorithms. For other networks, such as “FB Simmons83”, the values found by the algorithms are more similar and more Pareto optimal solutions are found, indicating many small improvements are made in uniqueness during the anonymization process.
For the different metrics, we observe two types of behavior. First, for the largest connected component size, no large difference is required for the best algorithms unless one wants to achieve small uniqueness values. For this task, many edge deletions can be made until a larger change occurs, which is also shown by the results in Appendix D. Especially if preserving the giant component is of high importance, it would be worth considering the partial variant of the problem.
Moving to community and betweenness centrality, we first observe that a large difference is lost in order to decrease uniqueness. This result can also be seen in Appendix D, where deleting the first percentage of edges results in a large difference in both the most central nodes and communities. After this decrease, the difference increases for both metrics, but grows less quickly for the community aspect.
Figure 5: Pareto optimal solutions in terms of uniqueness (horizontal axis) and performance on downstream tasks (vertical axis). Each dot represents a solution on the Pareto front found by an anonymization algorithm (color). Filled dots represent solutions for which less than 5% is deleted.
6.8 Runtime
Figure 6 shows the observed runtime of each algorithm, on each of the three variants, for each network. The runtime consists of computing probabilities to choose edges, for which the computational complexity varies for the anonymization algorithms introduced in Section 5.3, deleting the chosen edges from the network, which can be done in linear time, and accordingly determining the new anonymity of nodes by updating the equivalence partition (lines 6 and 7 in Algorithm 1). For this last operation, which is a computationally expensive part of the anonymization process, the runtime is affected by the number of affected nodes and the density of the network.
Comparing the different problem variants, the lowest runtimes are usually observed for the the budgeted variant, except for networks that do not require edge deletions for the partial variant if their initial uniqueness is less than 5%. Runtimes for the partial and full variant are often very similar. This small difference is likely due to the fact that later iterations require less runtime, which is further supported by results on the runtime per iteration in Appendix F.
Overall the runtime increases for networks with more nodes. At the same time, the first four networks show that a high uniqueness and density increase runtimes. The latter makes both measuring anonymity and computing the number of affected nodes more computationally expensive. For the four largest networks, which were included to assess the scalability of the algorithms, we observed runtimes of over 30 minutes for the budgeted variant for all algorithms, including es which spends a minimal amount of time on selecting edges.
Moving to the different algorithms, we first note that for many of the networks, there is no large difference in the runtimes required by the algorithms, implying that even on large networks the runtime can be high due to the bookkeeping operations. As shown in Appendix F, the time required for bookkeeping operations, including updating the equivalence classes, is usually much larger than the time used for the anonymization algorithms. Exceptions are networks such as “FB Simmons” for which the unique and ua spend the largest amount of time on the anonymization algorithm itself. However, for some networks, such as “Air traffic control” and “Network science”, the ua algorithm is slightly faster than the other algorithms, including es. This can be explained by the fact that ua is more effective in anonymization, which results in fewer required iterations and a lower runtime.
Figure 6: Runtimes for the five algorithms (color) with recompute gap $|E|/100$. Runtimes are shown for the three variants, full (bar), partial (circle) and budgeted (triangle). For the four largest networks, only runtimes of the budgeted variant are included.
7 Conclusion
In this paper, we introduced the anonymization problem in social networks in the form of three variants being 1) full, 2) partial and 3) budgeted anonymization. We introduced a generic algorithmic framework coined ANO-NET which allows one to anonymize networks, implementing the three variants, four measures for anonymity, the proposed anonymization algorithms and relevant parameters.
Of the edge alteration operations that can be applied, results on graph models showed that random edge deletion is most effective. Experiments to assess the effect of different anonymity measures on the anonymization process highlighted how networks are more difficult to anonymize using a more strict measure. The choice in measure affects both the initial uniqueness and how quickly uniqueness decreases. We proposed and experimentally evaluated five (of which four new) measure-agnostic anonymization algorithms. The effectiveness of the algorithms differed noticeably per network. In particular, networks with lower initial uniqueness and higher average path length showed to be easier to anonymize. From the anonymization algorithms one of the uniqueness based algorithms, ua, which accounts for the number of unique nodes affected, consistently showed to be most effective for anonymization. Compared to the baseline, this heuristic algorithm anonymized on average 4.8 times more nodes for the budgeted variant, and preserved 13.9 and 1.8 times more edges for the full and partial variant. Additionally, this algorithm managed to preserve three of the utility metrics better in budgeted anonymization. Furthermore, the ua algorithm appears to result in the best trade-off between anonymity and utility. While overall comparable, runtimes for the uniqueness based algorithms appear affected by the initial uniqueness and network density.
The aim of this paper was to lay the ground work for studying a general version of the anonymization problem and its variants, to study various properties of the problem, and evaluate five relatively simple heuristic algorithms both in terms of improving anonymity, and preserving data utility. An obvious next step is to design metaheuristic anonymization algorithms, or to explore the use of machine learning. One can also consider the use of the addition or rewiring operation, possibly in combination with deletion. To account for more attacker scenarios, it would be interesting to research which network properties, which may not be captured by the anonymity measures, could be used for de-anonymization, and design extensions accordingly. These extensions could both be based on structure, or account for additional network properties such as node or edge labels, or even temporality. Lastly, a step forward would be to dynamically account for how data utility decreases during the anonymization process and decide to continue or stop the anonymization process on-the-fly.
Appendix A Recompute gap
This appendix accompanies Section 5.2 of the full paper. Here, we compare the performance and runtimes achieved by the anonymization algorithms when using different recompute gaps. Figure 7 shows the results for different algorithms on a subset of the network datasets when using different recompute gaps. The recompute gap value of 1 implies that values are recomputed after each edge deletion, $|E|/100$ equals the default setting. The figure overall shows that using a smaller recompute gap similar results are obtained. Turning to the runtimes in Figure 8 we see that using a smaller recompute gap does result in larger runtimes.
Figure 7: Anonymization process for a selection of networks, anonymization algorithms and different recompute gaps (color).
Figure 8: Runtimes for anonymization using different recompute gap values (1, $|E|$/200 and $|E|$/100). Opaque values denote the largest average runtime of the anonymization algorithms, while the solid values denote the lowest value found.
Appendix B Anonymization process
This appendix contains supplementary figures accompanying Section 6.3. Results shows the full anonymization process, i.e., up to deleting all edges. Figure 9 shows results for all anonymity measures for all networks not shown in Figure 2 of the full paper.
Figure 9: Anonymization process for all networks, anonymization algorithms (line style) and measures (color).
Appendix C Properties and anonymization
This appendix accompanies Section 6.4 of the full paper and contains results on correlations between network properties and the effectiveness of the anonymization algorithms on the three variants of the anonymization problem. Table 3 denotes the Pearson correlation and p-value. For combinations deemed significant, Figure 10 shows the values for all networks.
1. Full 2. Partial 3. Budgeted
Network property
Pearson
correlation
p-value
Pearson
correlation
p-value
Pearson
correlation
p-value
Unique start -0.86 0.00 -0.92 0.00 -0.76 0.00
$|V|$ 0.13 0.62 0.29 0.25 0.06 0.81
Average degree -0.19 0.45 -0.13 0.61 -0.15 0.56
Median degree -0.80 0.00 -0.87 0.00 -0.68 0.00
Transitivity -0.52 0.03 -0.48 0.05 -0.59 0.01
Assortativity 0.15 0.56 0.29 0.25 0.10 0.71
diameter 0.64 0.01 0.60 0.01 0.62 0.01
Average path length 0.65 0.00 0.62 0.01 0.64 0.01
Table 3: Graph properties and correlation with anonymization results. Given various graph measures, denoted in the leftmost column, the Pearson correlation is shown with the best outcomes of the three anonymization variants (second column to last column). For each outcome both the Pearson correlation and $p$-value are given. Values with $p<0.05$ and Pearson correlation larger than 0.4, or smaller than -0.4, are shown in bold.
Figure 10: Network properties and anonymization. Each plot shows the the best value obtained (vertical axis) for each anonymization variant (color) on a specific network and the value of a network property (horizontal axis). Combinations plotted are those with a correlation deemed significant as denoted in Table 3.
Appendix D Deleting edges
This appendix accompanies Sections 6.6 of the full paper. Each figure shows how a network property or performance on a downstream task changes when deleting more edges, using the five anonymization algorithms.
Figure 11: Change in network properties when deleting edges. For each network (rows) and property (columns) the results show how data utility changes when deleting more edges (horizontal axis) using each of the five anonymization algorithms (color).
Appendix E Pareto fronts
This appendix accompanies Section 6.7 of the full paper. Figure 12 shows for each network the Pareto optimal solutions found by the algorithms, for each of the utility metrics.
Figure 12: Pareto fronts in terms of uniqueness (horizontal axis) and utility (vertical axis). For each network (rows), the results show for each property (columns) the Pareto optimal points found by each algorithm (color). Lines show one standard deviation.
Appendix F Runtimes
This appendix accompanies Section 6.8 of the full paper. Each figure shows both the full runtime, and time spent on the anonymization algorithm for a given network and anonymization algorithm.
Figure 13: Runtime per iteration. Each figure shows for each combination of network (rows) and algorithm (columns and color) the runtime per iteration. The color corresponding to the algorithm indicates the full runtime, the black dots the time spent on the anonymization algorithm only.